Adverse Impact Exercise

Symco Industries has been in a rapid growth mode for the past year. The company has worked hard to recruit the number of employees needed to fill organizational vacancies. However, because of the pace of recruitment and hiring, the company has not had an opportunity to look at its hiring data with regard to EEO considerations. A sequential, non-compensatory hiring approach is employed. That is, candidates must “pass” each stage to be hired. Below are the recruitment and hiring statistics for 2019 for the “professional” occupational category. Using the 4/5 rule as a guide, examine the data and answer the following questions:

(a) From the data presented below, is there any “overall” indication of discrimination? If yes, what type of discrimination is occurring? How do you know? Show all calculations.

(b) Examine the entire employment process. Is there potential for discrimination occurring at any stage in the process? If yes, who is being discriminated against and where in the process is discrimination occurring? Show all calculations.

(c) What argument, if any, could the company make in response to charges of discrimination based on the adverse impact analysis?

Employment Stage

Men

Women

Minority

Non-Minority

Applications Received

1200

400

800

800

Applicants Interviewed

400

100

250

250

Applicants Tested

200

50

50

200

Job Offers Made

120

20

20

120
Answer #1

(a) % of job offers made vs applications received of Non-minority = 120/800 = 15%

% of job offers made vs applications received of Minority = 20/800 = 2.5%

Critical Ratio = 2.5%/15% = 17%

This is less than 80%

Therefore, there is an Overall indication of discrimination

(b) There is potential discrimination in women at applicants tested stage, because % of applications tested vs applicants interviewed is 50/250  = 20%

(c) The company can make argument of bona fide occupational qualification (BFOQ) which requires more employees over females.

************Please please please rate the answer.It really helps em.Thanks!******************